OpenStudy (kj4uts):

How do you convert square root to decimals for example 5√2?

OpenStudy (kj4uts):

I know the answer is 7.1 rounded to one decimal place but my calculator has always figured it out for me and I would like to know how to solve it by hand.

OpenStudy (jdoe0001):

well... by hand can be a bit of unfinished work because \(\bf \sqrt{2}\) is an irrational number for one silly factoid, ancient greeks didn't like \(\sqrt{2}\) because it was irrational, among other irrational numbers =)

OpenStudy (mathmale):

In most cases you can't do an exact conversion of square roots to decimals. Instead, you'll have to use a calculator to find the desired root and then round that off to the desired or specified number of decimal places.

OpenStudy (jdoe0001):

I recall taking a square root manually when in school, just a simple quantity took about half a page =), and that was just the square root, not a cubic or quartic =) and if I recall correctly, the number was a close approximation, not exact per se

OpenStudy (kj4uts):

Oh the reason I am wondering is because I always used a converter online not my calculator (I accidentally said calculator) and I cannot use converter for my test. But am not sure how to solve in on my ti-84 when I plug it in it gives me it again 5 √2 again not in decimal form.

OpenStudy (jdoe0001):

hmmm I have hmm actually I have a ti-84 per se.. but mostly use the ti-83 plus....if I recall correctly, that's onthe [mode] or settings you have, there's also a [Math] button you can use, that has conversion options, to float, or fractions or decimals and such

OpenStudy (kj4uts):

ok I will check, so there are some settings for that?

OpenStudy (jdoe0001):

yes, usually... lemme turn it on

OpenStudy (jdoe0001):

do you happen to have an Android phone/tablet by any chance?

OpenStudy (kj4uts):

Never mind it works, I was originally trying in on my old ti-30 so I thought it did the same on the 84 but it converted it, sorry about that.

OpenStudy (jdoe0001):

hmm rats, I don't have a ti-84, just 86 and 82 and 83plus hmm but on the ti-83plus, you can do conversion from the [Math] button

OpenStudy (kj4uts):

Yeah I got it to work thanks for the help, I am reviewing my study guide do you know by any chance how to find the slope of a perpendicular line?

OpenStudy (jdoe0001):

a perpendicular line? perpendicular to what though? the Equator?

OpenStudy (kj4uts):

3x+12y=-18

OpenStudy (jdoe0001):

ok..... what's the slope of that line anyway?

OpenStudy (jdoe0001):

or better yet, can you solve that for "y"? what does it look like?

OpenStudy (kj4uts):

isn't the slope form y=mx+b so wouldent the 3 be the slope?

OpenStudy (mathmale):

Right. Take the given line, find its slope, and from that slope, find the slope of a line perpendicular to the given line.

OpenStudy (kj4uts):

so how do i solve for y do I have to get it to the other side

OpenStudy (jdoe0001):

yeap, the slope of that lin is 3 so, notice what mathmale said the NEGATIVE RECIPROCAL of that, is the slope of a line perpendicular to it thus \(\bf slope=3\implies \cfrac{3}{{\color{blue}{ 1}}}\qquad negative\implies -\cfrac{3}{{\color{blue}{ 1}}}\qquad reciprocal\implies - \cfrac{{\color{blue}{ 1}}}{3}\)

OpenStudy (kj4uts):

Ok I see now... so reciprocal is just flipping it

OpenStudy (jdoe0001):

pretty much, yes

OpenStudy (kj4uts):

now do I take that and plug it into the formula

OpenStudy (jdoe0001):

well... into the point-slope form, yes, assuming that you have two points to work with

OpenStudy (jdoe0001):

rather one point to work with, that is, an x,y pair

OpenStudy (kj4uts):

what do I do with the -1/3

OpenStudy (jdoe0001):

well, you can caress it, nurture it =) I mean.... what's the context? if you're looking for the equation of a line that is perpendicular to 3x+12y=-18 first off, you need a x,y pair and you can get that from that 3x=+12y=-18 pick any "x" value get the "y: use that x,y pair and then plug those two in the point-slope form, along with the negative reciprocal slope to get the equation of a line AT THAT POINT that is perpendicular to 3x+12y=-18 say... for example, you end up with x = who knows say 7, and y = 13 so then \(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 7}}\quad ,&{\color{blue}{ 13}})\quad % (c,d) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies -\cfrac{1}{3} \\ \quad \\ % point-slope intercept y-{\color{blue}{ 13}}={\color{green}{ -\frac{1}{3}}}(x-{\color{red}{ 7}})\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

3x=+12y=-18 meant 3ex+12y=-18, jus to clarify =)

OpenStudy (jdoe0001):

anyhow, so, assuming you get the pair 7,13 from that but you can simply pick any "x", get a "y", and use the slope and slap them together in the point-slope form :)

OpenStudy (jdoe0001):

rats... man, trying to clarify a typo, I end up with another typo =( 3x=+12y=-18 meant 3ex+12y=-18, jus to clarify =) I mean 3x+12y=-18 just to clarify the clarification =)

OpenStudy (jdoe0001):

anyhow... need to dash :)

OpenStudy (kj4uts):

ok thanks for your time and help I got 4